# Large Language Models Meet Knowledge Graphs for Question Answering: Synthesis and Opportunities

### v1

@@ 6 Conclusion
This survey has systematically examined the synthesis of LLMs and KGs in QA, presenting a novel taxonomy that categorizes methodologies based on QA types and the role of KGs. Our analysis highlights the strengths and limitations of current approaches, emphasizing the potential opportunities of leveraging the KGs to augment LLMs to overcome challenges such as limited reasoning capabilities, outdated knowledge, and hallucinations in complex QA scenarios. Despite significant advancements, several remaining challenges include efficient reasoning, knowledge alignment, and explainable QA. Future research should focus on developing adaptive frameworks that dynamically integrate up-to-date knowledge with LLMs, as well as establishing efficient methods for scaling reasoning, explainability, and fairness.
@@
This survey covers the taxonomy and summary of the recent advancements in synthesizing LLMs and KGs for QA and the challenges of complex QA, however, we are aware that this survey may miss some newly released works due to the rapid expansion of works on this topic. Moreover, the survey mainly highlights the alignments between the recent methodologies of incorporating LLMs and KGs for QA and the challenges of the various complex QA tasks, while these taxonomies from different perspectives are non-exclusive, and the overlap between the two taxonomies may arise. Furthermore, this survey underemphasizes the quantitative and experimental evaluation of different methodologies due to the various implementation details, the diversity of the benchmark datasets, and non-standardized evaluation metrics.
@@ A.1 Related Survey
Previous surveys have drawn the roadmap of unifying LLMs+KGs (Pan et al., 2024), have discussed the opportunities and challenges (Pan et al., 2023) of leveraging LLMs for knowledge extraction, ontologies and KGs constructions, have summarized the integration and synthesis paradigms of LLMs and KGs (Kau et al., 2024; Ibrahim et al., 2024), and gave an overview of the knowledge injection methods between LLMs and domain-specific knowledge (Song et al., 2025). Additionally, the existing surveys give an overview of multilingual knowledge graph question answering  (Perevalov et al., 2024), review temporal knowledge graph QA (Su et al., 2024), complex QA (Daull et al., 2023), and discuss search engines, KGs and LLMs from the perspective of user information seeking for QA (Hogan et al., 2025). The comparative analysis of these surveys across LLMs, KGs, LLMs+KGs, GraphRAG, and QA is given in Table 1.
@@ A.2 Paper Selection
We first retrieve research papers since 2021 by using Google Scholar and PaSa [^2] https://pasa-agent.ai/ using the phrases “knowledge graph and language model for question answering" and “KG and LLM for QA", and extend the search scope of the benchmark dataset paper to 2019, and then screen and select them based on their relevancy and publication venue quality.
@@ B.1 Complex QA
We divide the complex QA into the following categories according to the technical challenges.
@@
  • Multi-run and conversational QA. It involves user interactions to determine the final answer since some complex questions can not be answered in a single turn.
@@ B.2 Approaches of Synthesizing LLMs and KGs for QA
  • KG as Background Knowledge. When KGs are used as background knowledge to enhance LLMs for QA, the questions are parsed to identify the relevant subgraphs from KGs, then they are integrated with LLMs based on knowledge integration and GraphRAG.
@@ C.1 Summary Tables of Approaches
Table 3: Summary and Comparison of Methods in the Category of KG as Background Knowledge. NR: Newly-learned Rate, RR: Remembering Rate, CE Score: Cross Entropy Score, EM: Exact Match, MRR: Mean Reciprocal Rank, BLEU: Bilingual Evaluation Understudy, ROUGE: Recall-Oriented Understudy, NDCG: Normalized Discounted Cumulative Gain, METEOR: Metric for Evaluation of Translation with Explicit Ordering.
@@ C.2 Alignment of Approaches to Complex QA
Approach Multi-doc QA Multi-modal QA Multi-hop QA Multi-run QA XQA Temporal QA
@@ Appendix D Extended Optimization
Index-based Optimization. It is a time-consuming and complex task to create indexes from long-range facts and retrieve the relevant knowledge from large-scale graphs. To address this issue, GoR (Zhang et al., 2024b) leverages GNN and BERT score-based objectives to optimize node embeddings during graph indexing. KG-Retriever (Chen et al., 2024b) leverages a hierarchical index graph to enhance knowledge correlations and improve information retrieval for efficient knowledge indexing. NodeRAG (Xu et al., 2025b) integrates the heterogeneous graphs and fine-grained retrieval with RAG by optimizing the indexing of graph structures. DRO Shi et al. (2025) proposes a direct retrieval-agumentation method to directly estimate the distribution of document permutations from the select model for jointly learning the selection model and generative LM model.
@@
Prompt-based Optimization. To facilitate the deep fusion between the retrieved knowledge from KGs with the internal knowledge of LLMs, several prompt-based optimization approaches are proposed. For instance, StraGo (Wu et al., 2024c) enhances the quality and stability of the prompts based on the insights and strategic guidance learned from the historical prompts by using in-context learning. Meanwhile, several ranking strategies have been exploited to retrieve the top relevant knowledge from diverse knowledge bases, and the most relevant contexts are further fed to LLMs together with the prompts. For example, KG-Rank (Yang et al., 2024) leverages re-ranking techniques based on the score measuring relevance and redundancy to rank the top-$k$ triples from KGs and then combine them with the prompt to generate the answers for QA. Similarly, KS-LLM (Zheng et al., 2024b) introduces the evidence sentence selection module to retrieve the most relevant evidence sentences based on the ranking of the Euclidean distance between the triples and each evidence sentence. Instead of direct ranking of context-based relevance, BriefContext (Zhang et al., 2024a) introduces the preflight check to predict the relevance ranking between the user query and the retrieved documents, which is divided into multiple chunks for multiple RAG subtasks based on the map-reduce strategy. To leverage the multi-source knowledge for RAG-based QA, QUASAR (Christmann and Weikum, 2024) enhances RAG-based QA by effectively integrating unstructured text, structured tables, and knowledge graphs for evidence retrieval and re-ranking and filtering the relevant evidence from the retrieved evidence.
@@
Cost-based Optimization. To implement efficient and effective knowledge probing of LLMs, GLens (Zheng et al., 2024a) initially leverages the Thompson sampling strategy to measure the alignment between KGs and LLMs for addressing the knowledge blind spots of LLMs, and then designs a graph-guided question generator to convert KGs to text and designs a sampling strategy on the parameterized KG structure for accelerating KG traversal. Similarly, Coke (Dong et al., 2024a) minimizes the calls of LLMs for KGQA by introducing a cluster-level Thompson sampling to formulate the accuracy expectation and an optimized context-aware policy to distinguish the expert model based on question semantics. LMQL (Beurer-Kellner et al., 2023) minimizes the call of LLMs by generating an efficient inference procedure based on the LMP (Language Model Programming) constraints and control flow. It differs from the above work that CGPE (Tao et al., 2024) optimizes the knowledge retrieval based on clue-guided path exploration and information matching from knowledge bases to enhance the capabilities of LLMs for unfamiliar questions and reduce the costs of LLMs.
@@
To summarize, even if several optimizations and ranking strategies have been recently investigated to reduce the costs of graph retrieval, graph reasoning, and the length of the context of LLMs, however, the relevant subgraphs extraction, graph reasoning, and vector-based retrieval remain a computationally costly task.
@@ Appendix E Evaluations and Applications
The details of evaluation metrics, benchmark datasets, industrial and scientific applications, and demos of synthesizing LLMs and KGs for QA are summarized and compared as follows.
@@ E.1 Evaluations
Metrics. The evaluation metrics in synthesizing LLMs with KGs for QA are: (1) the metrics measuring the answer quality, BERTScore (Peng et al., 2024), answer relevance (AR), hallucination (HAL) Yang et al. (2025), accuracy matching, human-verficed completeness (Yu and McQuade, 2025); (2) the metrics measuring the retrieval quality of RAG, context relevance (Es et al., 2024), faithfulness score (FS) (Yang et al., 2024), precision, context recall (Yu et al., 2024; Huang et al., 2025), mean reciprocal rank (MRR)  (Xu et al., 2024), normalized discounted cumulative gain (NDCG) (Xu et al., 2024); (3) the metrics measuring the reasoning quality for multi-hop QA, Hop-Acc (Gu et al., 2024), reasoning accuracy (RA) (Li et al., 2025a).
@@
WebQSP (tau Yih et al., 2016) KBQA $\triangle$ Contains SPARQL queries for knowledge-based QA.
@@
CR-LT KGQA (Guo et al., 2024) KGQA $\triangle$ Long-tail entities and commonsense reasoning.
@@
CommonsenseQA (Talmor et al., 2019) Multi-choice QA $\triangle$ Contains questions with the correct answer and four distractor answers.
MedQA (Jin et al., 2021) Multi-choice QA $\triangle$ $\triangle$ Medical multi-choice QA dataset with multilingual medical examination text
@@
FanOutQA (Zhu et al., 2024) Multi-Hop QA $\triangle$ $\triangle$ Contains multi-hop question over multi-document.
@@
PATQA (Meem et al., 2024) Temporal QA $\triangle$ Present-anchored temporal QA
MenatQA (Wei et al., 2023) Temporal QA $\triangle$ $\triangle$ Evaluates the temporal reasoning capability of LLMs
@@
LLM-KG-Bench (Meyer et al., 2023) LLM-KG QA LLMs in knowledge graph engineering.
XplainLLM (Chen et al., 2024c) LLM-KG QA $\triangle$ Focuses on QA explainability and reasoning
@@
LiHua-World (Fan et al., 2025) LLM-KG QA Evaluates capability of LLMs on multi-hop QA in the scenario of RAG.
@@ E.2 Applications
We showcase the applications and demos in synthesizing LLMs with KGs for QA.
KAG (Liang et al., 2024a) (by Antgroup) [^3] https://github.com/OpenSPG/KAG is a domain-knowledge augmented generation framework that leverages KGs and vector retrieval to bi-directionally enhance LLMs for knowledge-intensive tasks, such as QA.
@@
GraphRAG-QA (by NebulaGraph) [^5] https://github.com/wey-gu/demo-kg-build is an industrial demo of GraphRAG integrating several query engines for augmenting QA, NLP2Cypher-based KG query engine Wu et al. (2022), vector RAG query engine, and Graph vector RAG query engine.
@@
Fact Finder Steinigen et al. (2024) (by Fraunhofer IAIS and Bayer) [^7] https://github.com/chrschy/fact-finder is a QA system that augments LLMs with the query-based KG retrieval from medical KG to improve the correctness and completeness of the answers.
AprèsCoT (Shirdel et al., 2025) (by University of Waterloo) [^8] http://lg-research-2.uwaterloo.ca:8050/aprescot is a visualization tool for understanding and explaining the answers generated by LLMs with CoT-based prompting and visualized path exploration and inference over KGs.

### v2

@@ 6 Conclusion
This survey has systematically examined the synthesis of LLMs and KGs in question answering, presenting a novel taxonomy that categorizes methodologies based on QA types and the role of KGs. Our analysis highlights the strengths and limitations of current approaches, emphasizing the potential opportunities of leveraging the KGs to augment LLMs to overcome challenges such as hallucinations, limited reasoning capabilities, and knowledge conflicts in complex QA scenarios. Despite significant advancements, several remaining challenges include efficient knowledge retrieval, dynamic knowledge integration, effective reasoning over knowledge at scale, and explainable and fairness-aware QA. Future research should focus on developing adaptive frameworks that dynamically integrate up-to-date knowledge with LLMs, as well as establishing efficient methods for scaling reasoning, explainability, and fairness.
@@
This survey covers the taxonomy and survey of the recent advancements in synthesizing LLMs and KGs for QA, as well as the discussion of its challenges and opportunities. However, we are aware that this survey may miss some newly released works due to the rapid expansion of works on this topic. Moreover, the survey mainly highlights the alignments between the recent methodologies of incorporating LLMs and KGs for QA and the challenges of the various complex QA tasks, while these taxonomies from different perspectives are non-exclusive, and the overlap between the two taxonomies may arise. Furthermore, this survey underemphasizes the quantitative and experimental evaluation of different methodologies due to the various implementation details, the diversity of the benchmark datasets, and non-standardized evaluation metrics.
Acknowledgments
This work was supported by the Novo Nordisk Foundation grant (NNF22OC0072415), the National Natural Science Foundation of China (U23B2057, 62176185, 62376058), and the GOBLIN COST Action (CA23147).
@@ A.1 Related Survey
Previous surveys have drawn the roadmap of unifying LLMs+KGs (Pan et al., 2024), discussed the opportunities and challenges (Pan et al., 2023) of leveraging LLMs for knowledge extraction, ontologies and KGs constructions, summarized the integration and synthesis paradigms of LLMs and KGs (Kau et al., 2024; Ibrahim et al., 2024), and gave an overview of the knowledge injection methods between LLMs and domain-specific knowledge (Song et al., 2025). Additionally, the existing surveys give an overview of multilingual knowledge graph question answering  (Perevalov et al., 2024), review temporal knowledge graph QA (Su et al., 2024), complex QA (Daull et al., 2023), and discuss search engines, KGs and LLMs from the perspective of user information seeking for QA (Hogan et al., 2025). The comparative analysis of these surveys across LLMs, KGs, LLMs+KGs, GraphRAG, and QA is given in Table 1.
@@ A.2 Paper Selection
We first retrieve research papers since 2021 by using Google Scholar and PaSa [^2] https://pasa-agent.ai/ using the phrases “knowledge graph and language model for question answering" and “KG and LLM for QA", and extend the search scope of the benchmark dataset paper to 2016, and then screen and select them based on their relevancy and publication venue quality.
@@ B.1 Complex QA
We classify the complex QA into the following categories according to the technical challenges.
@@
  • Conversational QA. It involves user engagement in multi-turn QA to understand the given context, determine the final answers, and satisfy their information needs.
@@ B.2 Approaches of Synthesizing LLMs and KGs for QA
  • KG as Background Knowledge. When KGs are used as background knowledge to enhance LLMs for QA, the questions are parsed to identify the relevant subgraphs from KGs, which are then incorporated with LLMs based on knowledge integration and fusion.
@@ C.1 Summary Tables of Approaches
Table 3: Summary and Comparison of Methods in the Category of KG as Background Knowledge. NR: Newly-learned Rate, $R$: Remembering Rate, CE Score: Cross Entropy Score, EM: Exact Match, MRR: Mean Reciprocal Rank, BLEU: Bilingual Evaluation Understudy, ROUGE: Recall-Oriented Understudy, NDCG: Normalized Discounted Cumulative Gain, METEOR: Metric for Evaluation of Translation with Explicit Ordering.
@@ C.2 Alignment of Approaches to Complex QA
Approach Multi-doc QA Multi-modal QA Multi-hop QA Conversatiional QA XQA Temporal QA
@@ Appendix D Extended Optimization
Index-based Optimization. It is a time-consuming and complex task to create vector indexes from long-range facts and retrieve the relevant knowledge from large-scale graphs. To address this issue, GoR (Zhang et al., 2024b) leverages GNN and BERT score-based objectives to optimize node embeddings during graph indexing. KG-Retriever (Chen et al., 2024c) leverages a hierarchical index graph to enhance knowledge correlations and improve information retrieval for efficient knowledge indexing. NodeRAG (Xu et al., 2025b) integrates the heterogeneous graphs and fine-grained retrieval with RAG by optimizing the indexing of graph structures. DRO (Shi et al., 2025) proposes a direct retrieval-agumentation method to directly estimate the distribution of document permutations from the selected model for jointly learning the selection model and generative LM model.
@@
Prompt-based Optimization. To facilitate the deep fusion between the retrieved knowledge from KGs and the internal knowledge of LLMs, several prompt-based optimization approaches are proposed. For instance, StraGo (Wu et al., 2024c) enhances the quality and stability of the prompts based on the insights and strategic guidance learned from the historical prompts by using in-context learning. Meanwhile, several ranking strategies have been exploited to retrieve the top relevant knowledge from diverse knowledge bases, and the most relevant contexts are further fed to LLMs together with the prompts. For example, KG-Rank (Yang et al., 2024) leverages re-ranking techniques based on the score measuring relevance and redundancy to rank the top-$k$ triples from KGs and then combine them with the prompt to generate the answers for QA. Similarly, KS-LLM (Zheng et al., 2024b) introduces the evidence sentence selection module to retrieve the most relevant evidence sentences based on the ranking of the Euclidean distance between the triples and each evidence sentence. Instead of direct ranking of context-based relevance, BriefContext (Zhang et al., 2024a) introduces the preflight check to predict the relevance ranking between the user query and the retrieved documents, which is divided into multiple chunks for multiple RAG subtasks based on the map-reduce strategy. To leverage the multi-source knowledge for RAG-based QA, QUASAR (Christmann and Weikum, 2024) enhances RAG-based QA by effectively integrating unstructured text, structured tables, and knowledge graphs for evidence retrieval and re-ranking and filtering the relevant evidence from the retrieved evidence.
@@
Cost-based Optimization. To implement efficient and effective knowledge probing of LLMs, GLens (Zheng et al., 2024a) initially leverages the Thompson sampling strategy to measure the alignment between KGs and LLMs for addressing the knowledge blind spots of LLMs, and then designs a graph-guided question generator to convert KGs to text, together with a sampling strategy on the parameterized KG structure to accelerate KG traversal. Similarly, Coke (Dong et al., 2024a) minimizes the calls of LLMs for KGQA by introducing a cluster-level Thompson sampling to formulate the accuracy expectation and an optimized context-aware policy to distinguish the expert model based on question semantics. LMQL (Beurer-Kellner et al., 2023) minimizes the call of LLMs by generating an efficient inference procedure based on the LMP (Language Model Programming) constraints and control flow. It differs from the above work in that CGPE (Tao et al., 2024) optimizes the knowledge retrieval based on clue-guided path exploration and information matching from knowledge bases to enhance the capabilities of LLMs for unfamiliar questions and reduce the costs of LLMs.
@@
To summarize, even if several optimizations and ranking strategies have been recently investigated to reduce the costs of graph retrieval, graph reasoning, and the length of the context of LLMs. However, the relevant subgraphs extraction, graph reasoning, and vector-based retrieval remain a computationally costly task.
@@ Appendix E Evaluations and Applications
The details of evaluation metrics, benchmark datasets, industrial and scientific applications, and demonstrations of synthesizing LLMs and KGs for QA are summarized and compared as follows.
@@ E.1 Evaluations
Metrics. The evaluation metrics in synthesizing LLMs with KGs for QA are: (1) the metrics measuring the answer quality, BERTScore (Peng et al., 2024), answer relevance (AR), hallucination (HAL) (Yang et al., 2025), accuracy matching, human-verficed completeness (Yu and McQuade, 2025); (2) the metrics measuring the retrieval quality of RAG, context relevance (Es et al., 2024), faithfulness score (FS) (Yang et al., 2024), precision, context recall (Yu et al., 2024; Huang et al., 2025), mean reciprocal rank (MRR)  (Xu et al., 2024), normalized discounted cumulative gain (NDCG) (Xu et al., 2024); (3) the metrics measuring the reasoning quality for multi-hop QA, Hop-Acc (Gu et al., 2024), reasoning accuracy (RA) (Li et al., 2025a).
@@
WebQSP (tau Yih et al., 2016) KBQA $\triangle$ Includes SPARQL queries for knowledge-based QA.
@@
CR-LT KGQA (Guo et al., 2024) KGQA $\triangle$ Supports long-tail entities and commonsense reasoning.
@@
CommonsenseQA (Talmor et al., 2019) Multi-choice QA $\triangle$ Includes questions with the correct answer and four distractor answers.
MedQA (Jin et al., 2021) Multi-choice QA $\triangle$ $\triangle$ Medical multi-choice QA dataset with multilingual medical examination text.
@@
FanOutQA (Zhu et al., 2024) Multi-Hop QA $\triangle$ $\triangle$ Includes multi-hop question over multi-document.
@@
PATQA (Meem et al., 2024) Temporal QA $\triangle$ Provides Wikidata questions for Present-anchored temporal QA.
MenatQA (Wei et al., 2023) Temporal QA $\triangle$ $\triangle$ Evaluates the temporal reasoning capability of LLMs.
@@
LLM-KG-Bench (Meyer et al., 2023) LLM-KG QA Evaluates capabilities of LLMs in knowledge graph engineering.
XplainLLM (Chen et al., 2024d) LLM-KG QA $\triangle$ Focuses on QA explainability and reasoning.
@@
LiHua-World (Fan et al., 2025) LLM-KG QA Evaluates the capability of LLMs on multi-hop QA in the scenario of RAG.
@@
BlendQA (Xin et al., 2025) LLM-KG QA $\triangle$ Evaluates cross-knowledge source reasoning capabilities of RAG for QA.
@@ E.2 Applications
We showcase the applications and demonstrations in synthesizing LLMs with KGs for QA.
KAG (Liang et al., 2024a) (by Antgroup) [^3] https://github.com/OpenSPG/KAG is a domain-knowledge augmented generation framework that leverages KGs and vector retrieval to bidirectionally enhance LLMs for knowledge-intensive tasks, such as QA.
@@
GraphRAG-QA (by NebulaGraph) [^5] https://github.com/wey-gu/demo-kg-build is an industrial demo of GraphRAG integrating several query engines for augmenting QA, NLP2Cypher-based KG query engine (Wu et al., 2022), vector RAG query engine, and Graph vector RAG query engine.
@@
Fact Finder (Steinigen et al., 2024) (by Fraunhofer IAIS and Bayer) [^7] https://github.com/chrschy/fact-finder augments LLMs with the query-based retrieval from medical KG to improve the completeness and correctness of answers.
AprèsCoT (Shirdel et al., 2025) (by University of Waterloo) is a visualization tool for explaining the answers generated by LLMs with CoT-based path exploration and inference over KGs.